FEL Upgrade Project Weekly Brief
December 10-14, 2001

We are pleased to thank Governor Gilmore and his administration who announced this week a new $1B bond
issue for support of higher education in Virginia that includes a $3M line item for the planned addition to the
FEL User Facility building to house the compact synchrotron and injector test area. If the bond issue is
approved by the Virginia legislature in the upcoming Jan.-Mar. 2002 session and by the Virginia voters in
November 2002, then the funding would be available in July of 2003.  This same bond issue includes support
for many science programs (particularly materials science and bioscience) at Virginia’s research universities
that are involved with the FEL user program.

Jefferson Lab will have a DOE review of our nuclear physics operations budget next month on Jan. 22-24. The
FEL project management will be presenting the impact of the FEL project on the lab operations in terms of
contributions to maintaining the lab core competencies (SRF and other accelerator physics) and the supporting
labor for these efforts.

DOE approved the UV FEL project subcontract to the Aerospace Corporation for their effort toward the
design and fabrication of a UV laser microfabrication tool for use with the FEL Upgrade.  Our thanks to our PI
at Aerospace Corp, Henry Helvajian, our support from JLab business services (R. Lusk and T. Danforth) and
our DOE contracting officer, Wayne Skinner, for their assistance in this effort.

We are continuing discussions with Cornell University, which is working on a prototype energy recovered linac
x-ray source involving a source current of 100 mA , and with Daresbury Laboratory in the UK , which is also
working on a next generation light source, about potential collaborations.

Jefferson Lab has a tentative visit next Wednesday (Dec. 19th)by several Congressional staff members from
the Senate Armed Services Committee for a tour and briefing on the FEL program.

WBS 4 (Injector):
We completed a high voltage emission test on a stainless steel electrode that was implanted
in the re-commissioned plasma/ion implantation system operated by CWM. We are using
this test sample and others to benchmark the system before we implant the new electrodes for the Upgrade
gun. This test electrode was quiet up to 25 MV/m but started emitting above that level and would not condition
out.  The ball cathode was final machined (ready for polishing) and had the stock (wafer holder)
match-machined to the interior step.  The gun stand drawings were signed off and submitted for fabrication
bids. The wafer shield rear operation link and bellows drawings were signed off and submitted to the shop for

Gun HVPS Tank - Meetings were held this week to help firm-up the tank design.  No date has been set for the
release of the drawing package.  RCD was contacted this week concerning the combined running and
conditioning resistors.

WBS 6 (RF):
Injector Klystrons - The wiring of the 1st klystron continues.  The LCW water leaks have been stopped.  The
initial 1/2 power tests appear to be firming up for the 1st or 2nd week of January.  CPI was notified and they
may want to attend to witness these tests.

Injector HVPS - An outside vendor has been contacted to remove the present heavy components from the
HVPS's and install the new heavy components.  Their price is less than it would cost us to purchase the
necessary equipment.  The HVPS tear-out is starting Friday on Unit 3.  Unit 4 tear-out will start after the
klystron tests are completed.

WBS 8 (Instrumentation):
A trip was made to Varian this week to work out response time for their cold cathode controller. The demo
unit they had provided for upgrade work and SNS interlocks had been setup improperly.  The response time
has gone from 8 seconds to 2 millisecond.  Tune into the website for more details:  http://laser.jlab.org/sns.

An inventory control system is being developed to support the barcodes that were assigned to components prior
to commencing the rip-out portion of the upgrade.  This system is nearing completion and will be expanded to
include instrumentation chassis, test equipment and other "portable" equipment as requested.

The FLOG system is being enhanced to make it easier to access and provide a search feature.

Design of the new stepper motor control chassis is underway.  The intent is to double the number of channels
available without any increased "real estate" requirement.

Cable removal in the vault continues.

WBS 9 (Transport):
Optical Chicane Dipoles (GW)
o  Magnet Enterprises International in Oakland, CA, was under a cure notice on our two contracts, QX
    coils and GW complete magnets.  They spent the week concentrating on QX coils because of our
    request that they have higher priority.  Work slowed on the GW coil as they also resolved coil form
    dimensions.  The actual potting was delayed again to next week. In the interim, they started rough
    machining the core pieces.
Injector Dipoles (DU/DV)
o  Concerns about the extent of the injector dipole field integral quality in the return beam zone and
    about the pinch in the injector chamber forced by the field clam slider for the wide portion of the
    dipole.  In both cases there seems to be no problem.  DULY Research magnetic model results
    indicated the GV Dipoles have specification good field integral to the edge of the field clamp
    sliders. Further analysis indicates that we can also slide the above slider an additional 3/4 inch with
    no loss in field quality.
o  With the above results and the review of drawings and specifications we did this week, we should
    be able to sign off the drawings and get these magnets into procurement.  The requisition is signed
    and we will do another electronic procurement as we did successfully on the 180 degree (GY)
Arc Dipoles (GY, GX, GQ)
o  Advanced Energy Systems (AES) is revising the drawings of the pole tips of the Bend Dipole (GX)
    to reflect the magnetic model results.
o  Tom Schultheiss of AES is now bringing the Reverse Bend Dipole (GQ) magnetic model in
    conformance with the improvements to the GX
o  The responses to our questions on the 180° Dipoles (DY) were received on Tuesday and
    evaluated.  We are requesting Best And Final Offers (BAFOs) which are due Monday.
Magnet Measurement
o  The parts for the dipole measurement stand are being fabricated.
o  Measurement of the corrector dipoles continued.
3 inch quad (QX)
o  The core vender, New England Techni-Coil (NTC) is waiting on our resolution of the quality of QX
    coils before assembling the first article.
o  Jeffrey Hubrig of Innovation Services continued monitoring the production of QX Coils at Magnet
    Enterprises International (MEI), bringing their cure temperature of the coils within epoxy cure
    guidelines by doing a calibration on the oven in a dry run followed by an actual cure per the lessons
    learned in the dry run.  Simultaneously, he reports a complete turnaround of the vendor on all issues
    raised in the cure notice of last week.  They should be able to continue production soon as directed
    by our Procurement Department.
o  The remaining issue of the lowered cure temperature used by MEI on the initial batches of coils
    was resolved.  Jeffrey Hubrig was able to test the epoxy hardness of a series of rejected coils from
    all the lots.  They all registered Shore hardness in the 80s while the maximum hardness reported by
    the epoxy manufacturer is 92 for an ideal cure.  These strengths indicate that though the cure is not
    ideal, the strength is substantial enough the coils will not be falling apart or warp during service.  We
    will therefor accept the coils and authorize NTC to start assembly.
Trim Quad (QT)
o  The magnets are in fabrication at Milhous Control of Virginia.
Sextupole (SF)
o  DULY Research continues to model the sextupole.
Octupole (OT)
o  Work on this magnet is on furlough until the sextupole is designed.
Corrector Dipoles (DB, DJ)
o  The magnets are being qualified on the measurement stand.
Beam Line and Vacuum
o  Work continues at AES on the design the Arc Chambers.
o  Design work on girders and stands continued. Parts for the regions between cryomodules are being
    prepared for signature.
o  The layout of the telescope in front of the first arc is complete, accommodating David Douglas’
    latest revision that works better for the UV line.
o  All the new three leg pedestals in fabrication.
o  Chuck Hutton has drawn up the brackets for the skew sextupoles that we will need at the ends of
    the Cryomodules.  He is now working on making a two pair of Haimson like Correctors, suitable for
    high temperature bake that have to be mounted in the gun solenoid.  (Haimson Research declined to
    make a high temperature version of their coils.)

WBS 10 (Wiggler):
Commissioning measurements were completed on wiggler 1.  Due to the need to divert resources to other
tasks, the commissioning measurements on wiggler 2 will be delayed a few weeks.  Analysis of the wiggler 1
measurements is ongoing.  Calculations of the spectrum from the entire wiggler using measured fields are
being used to figure out how to set up the wiggler for optimum lasing.  The setup for production of X-rays is
also being studied as a test case of the tunability.  It works differently than expected but it does work.
Continuous tunability of the wavelength is possible by keeping the relationship between the wiggler and
dispersion strengths fixed on a curve which has been derived.  The vacuum chamber is complete and awaiting
completion of final commissioning measurements for installation.

WBS 11 (Optics):
Progress for the upgrade was made in the following areas:

The ultraviewer was tested and found to meet specifications.  Several parts omissions were found in the
drawings, and revisions will be issued.  One technical issue (fairly easily solved) will be discussed before the
rest of the ultraviewers are assembled.

The detailing of the design for the optical cavities has reached the point where the drawing package for
components from the "big C" (the yaw axis yoke) through the mirror holders is now in review.  We received
the rotary feedthroughs and stepper motors for the optical cavity assembly y-axis motion.  The ball bearings for
the leadscrew have been cleaned for high vacuum use.  They appear to be usable as-is.  Because of the low
cost and potentially high payoff, one pair is going to be coated with dicronite and then evaluated for particulate
generation.  Purchase of the rails and pillow blocks for the optical cavities (plus spares for the test stand) are in

The OCMMS tests continue.  The new position sensitive detector arrived.

Decommissioning/Other Activities:
We identified parts of the IR Demo that we won't use for some time and moved them to the Metrology lab for
storage.  We assisted the Hall A Compton polarimeter team with the change out of the mirrors in the high
finesse cavity.  We gave a tour to David Sutter and Bruce Strauss of DOE HEP.